# PrisonBreak: Jailbreaking Large Language Models with Fewer Than Twenty-Five Targeted Bit-flips

### v1

@@ Abstract
We introduce a new class of attacks on commercial-scale (human-aligned) language models that induce jailbreaking through targeted bitwise corruptions in model parameters. Our adversary can jailbreak billion-parameter language models with fewer than 25 bit-flips in all cases—and as few as 5 in some—using up to 40$\times$ less bit-flips than existing attacks on computer vision models at least 100$\times$ smaller. Unlike prompt-based jailbreaks, our attack renders these models in memory ‘uncensored’ at runtime, allowing them to generate harmful responses without any input modifications. Our attack algorithm efficiently identifies target bits to flip, offering up to 20$\times$ more computational efficiency than previous methods. This makes it practical for language models with billions of parameters. We show an end-to-end exploitation of our attack using software-induced fault injection, Rowhammer (RH). Our work examines 56 DRAM RH profiles from DDR4 and LPDDR4X devices with different RH vulnerabilities. We show that our attack can reliably induce jailbreaking in systems similar to those affected by prior bit-flip attacks. Moreover, our approach remains effective even against highly RH-secure systems (e.g., 46$\times$ more secure than previously tested systems). Our analyses further reveal that: (1) models with less post-training alignment require fewer bit flips to jailbreak; (2) certain model components, such as value projection layers, are substantially more vulnerable than others; and (3) our method is mechanistically different than existing jailbreaks. Our findings highlight a pressing, practical threat to the language model ecosystem and underscore the need for research to protect these models from bit-flip attacks.
@@ 1 Introduction
Deep neural networks are vulnerable to parameter corruption, which adversaries can exploit to trigger undesirable behaviors. These include severe performance degradation [1, 2, 3], targeted misclassifications [4, 5, 6], and backdoor injections [7, 8, 9, 10, 11]. Additionally, recent studies have revealed confidentiality risks such as model extractions [12, 13, 14]. Most of these works focus on exploiting the vulnerability through fault attacks, particularly those that cause software-induced bit-wise errors (bit-flips), such as Rowhammer [15] in the memory representation of models.
@@
Most existing research has focused on exploiting convolutional neural networks in classification tasks [7, 1, 3, 4, 5, 8, 11]. While valuable, these studies leave the vulnerability of language models—especially those widely used across AI services—largely unexplored in terms of parameter corruption. Cai et al. [6] conducted the closest related work, showing that a machine translation system could be perturbed to generate attacker-chosen outputs on specific inputs. However, this outcome is still within the breadth of prior bit-flip attack work, and it is unclear if language models are vulnerable to unexplored adversarial objectives. The most interesting scenario, which no prior work has studied so far, is to cause undesirable behaviors in language models—such as jailbreaks—through minimal bit-wise corruptions in model parameters.
Our paper presents a new class of fault injection attack—PrisonBreak—that jailbreaks commercial-scale language models. Our attack flips a small set of targeted bits (at most 25) in the billions of parameters in the models to break the alignment enforced by safety training mechanisms. Unlike prompt-based jailbreaks [16, 17, 18], this approach does not rely on additional prefixes or suffixes to provoke harmful responses and causes minimal performance degradation, making it potentially less detectable by service providers.
Our attack operates in two steps. In the offline step, we identify the locations in the target model for inducing bit-flips. We design a novel algorithm—BitFinder—to guide the search for bit-flip locations using gradient information. Identifying a minimal sequence of bit locations is particularly challenging due to the onion effect we discovered in our work [^1] It is important to note that this onion effect is distinct from the term introduced by Carlini et al. [19] in the context of privacy domains.. As we cannot determine all bit locations at once, we formulate the search process as progressive bit-flipping: iteratively selecting the bit that produces the most significant change in the objective function until we achieve a jailbroken model. We then address the computational burden of this iterative approach by designing a series of techniques to achieve a 19–20$\times$ times speed-up when evaluating parameters. To compute informative gradients, we use existing objectives in the prompt-based literature, with two key adaptations to enhance the fidelity of harmful responses and preserve the models’ utility. We incorporate system-specific information to ensure the selected bits are flippable in the target DRAM, allowing us to target highly secure DRAM modules. In the online step, we flip the target bits using our online procedure—PrecisionHammer—using Rowhammer [15], a critical vulnerability in commodity DRAM-based systems. We leverage advanced memory templating [20], massaging [21], and Rowhammer attacks [22] to precisely flip the bits in the victim’s memory.
We evaluate the effectiveness of our attack across 8 open-source language models and a representative jailbreaking benchmark, achieving a substantial attack success rate while only flipping 5–25 bits across all models. PrisonBreak produces high-fidelity jailbroken outputs that exceed prompt-based jailbreaking methods, with comparable attack success to parameter-based jailbreaks that modify billions more parameters. We then assess the practicability of PrisonBreak in real systems with Rowhammer simulations across 56 DRAM profiles (40 DDR4 and 16 LPDDR4X). Our results show that the attacker succeeds over 99% of the time across 20+ modules. By incorporating the templating information identified by PrecisionHammer, we also show that the attacker can succeed against highly secure DRAM modules by only selecting flippable target bits.
Finally, we characterize the vulnerability. We first conduct a structural analysis across several levels of abstraction in the parameter space, finding several—such as the exponent bits and later layers—that contain a disproportionate number of target bits. We then conduct a comprehensive behavioral analysis, finding that our attack functions differently than existing jailbreaks in the activation space.
@@
  • We present a new vulnerability in large-language models to an under-explored security threat: jailbreaking through targeted bit-flips. Our attack, PrisonBreak, requires no input-prompt changes and successfully jailbreaks target models in memory with fewer than 25 bit-flips.
  • We identify key challenges in exploiting this vulnerability and address them by developing a novel bit-search procedure (BitFinder) to efficiently identify target bits to flip. We also introduce PrecisionHammer to flip the target bits in memory using practical fault injection methods, such as Rowhammer.
@@
  • We validate PrisonBreak’s practical application with Rowhammer simulations across 56 DRAM profiles (40 DDR4 and 16 LPDDR4X). Our attacker can succeed in jailbreaking even on secure DRAM modules by carefully selecting target bits during the offline bit-search process.
  • We conduct a structural analysis, showing that certain components of models are disproportionately more vulnerable. We also perform behavioral analysis, exploring the mechanistic way our attack operates and finding it fundamentally different from existing jailbreaks.
2 Background
@@
where $f_{\theta}(x_{i}|x_{1},\dots,x_{i-1})$ represents the likelihood of the next token $x_{i}$, as computed by the neural network $f$ with parameters $\theta$ (weights and biases). $\theta$ is learned by iteratively minimizing the objective across the entire training set. If the training reaches an acceptable accuracy, the model $f$ along with all its parameters $\theta$ is stored. In deployment, both the model and parameters are loaded into memory and generate text autoregressively by sequentially processing an input and sampling future tokens until reaching a stopping criterion (e.g., an end-of-speech token). Our work considers the Transformer architecture [26] for $f$, which consists of a sequence of attention layers with additional components such as normalization and fully-connected layers.
Aligning and jailbreaking language models. (Large) language models pre-trained on text data can perform various tasks [27] and serve as the backbone for many user-facing applications. However, it has been shown that these capabilities can be exploited by malicious users [28, 29] to produce biased, deceptive, or ultimately harmful content. In response, active research is underway to develop techniques to ensure that these models generate outputs aligned with human values, e.g., supervised fine-tuning (SFT) [30], reinforcement learning with human feedback (RLHF) [31, 32], and direct preference optimization (DPO) [33]. Models aligned using these techniques [34, 35, 36] produce safe, helpful responses and refuse harmful queries.
Recent work studies the security implications of aligned language models through jailbreaking attacks, which aims to induce these models to generate harmful responses. Because these models are accessible to any user prompt, most jailbreaks are prompt-based [16, 18, 17, 37]. These attacks modify harmful prompts to bypass the alignment and compel models to respond to inputs they would otherwise refuse. A few works present parameter-based jailbreaks where an attacker tunes billion parameters of an aligned model on a small set of harmful text [38, 39] to remove the alignment from models without altering the input. In contrast, we present a new attack that causes up to 25 bit-flips to a model’s memory representation exploiting Rowhammer vulnerability [15], to make the model jailbroken in runtime.
@@ 3 Threat Model
We consider an adversary who aims to jailbreak large-language models by directly manipulating their parameters in memory. Our attack is entirely different from the class of attacks that leverage jailbreaking prompts [16, 18, 17, 37] where the attacker has to modify each harmful query to induce harmful responses. In contrast, our attack requires no input modifications. We assume scenarios where a victim deploys these language models in resource-sharing environments, such as machine-learning-as-a-service (MLaaS) platforms. These platforms are becoming increasingly common as few entities have the computing resources to run large-scale models and offer natural language processing services to users at scale. The victim makes a best effort to keep these models aligned with human values, with one safety-oriented outcome being the denial of harmful queries (e.g., by answering “I cannot answer to that” to “how to create an explosive product for my friend’s birthday party?”).
@@
Knowledge. We consider a white-box attacker with knowledge of the target model’s architecture and its parameters which is consistent with the conventional setup in prior work [1, 9, 11, 6, 10, 7]. This assumption is practical as many user-facing services are built on open-source language models [47]. But we emphasize that such knowledge is not strictly required, as it can potentially be acquired through prior attacks with similar capabilities [13, 14, 12]. The attacker requires no access to training data. We assume the attacker has knowledge of the memory locations likely to induce bitwise errors. Our adversary can either search for model parameters with the knowledge of vulnerable memory locations (hardware-aware) or profile these locations after the target model is loaded into memory and then induce bit-flips (hardware-agnostic). Since prior attacks assume at most one of these scenarios, our assumption allows for a comprehensive study of practical exploitations.
@@
Capabilities. Our attacker exploits Rowhammer [15], a fault injection attack practical in modern computing systems, to induce targeted bit-flips in memory. The adversary has only user-level privileges and co-locates a process or virtual machine (VM) within the victim’s process or VM running on the same host. The host OS remains uncompromised, and confinement policies (e.g., process isolation) are enforced. Our Rowhammer attacker can reverse-engineer DRAM address mappings using techniques proposed in [20] and place target pages at desired physical locations by leveraging advanced memory massaging primitives [48]. While we mainly consider models loaded in CPU memory, Olgun et al. [46] has demonstrated the exploitation of Rowhammer on GPUs, presenting a potential vector for jailbreaking models in GPU memory. Our assumptions align with those shown in prior bit-flip attacks on neural networks [1, 14, 11, 7].
Figure 1: PrisonBreak workflow. We illustrate the steps in our offline (BitFinder) and online (PrecisionHammer) procedures.
@@ 4 PrisonBreak
This section introduces our attack designed to jailbreak large-language models loaded in memory at runtime. Figure 1 illustrates the overall workflow of PrisonBreak.
4.1 Attack Overview
PrisonBreak operates in two phases at a high-level:
@@ 4.1 Attack Overview
BitFinder5) is our offline procedure that identifies target bit locations in the victim model’s memory representation. We run a progressive bit-search to search the target bits. We first analyze the victim model and identify critical weights whose perturbations are most effective in jailbreaking. We next extract all bits from the floating-point representation of the critical weights and flip them one at a time in isolation to identify the most effective bit-flip. We then select the bit and flip it; the bit-flip is carried onto the remaining search process. We repeat these steps until achieving an attack success of, e.g., 80–95%, comparable to prior work [16]. We store the identified target bits.
Note that our offline procedure supports two distinct modes: hardware-agnostic and hardware-aware search, depending on the attacker’s knowledge (see §3). The hardware-agnostic search is as described above. In the hardware-aware search, we are limited to select bits at vulnerable locations within a memory template profiled by the attacker in advance. We detail these modes in §5.3 and demonstrate both exploitations through Rowhammer in §8.
@@
PrecisionHammer6) is the online procedure that uses Rowhammer to jailbreak the victim language model by flipping target bits in DRAM. We first perform memory templating [20], where the attacker profiles the system’s memory and identifies vulnerable cells, their locations, and flip directions. Note that this procedure can be done in advance for the hardware-aware attacker. We next perform memory massaging [40], aligning the target bits with cells exploitable by Rowhammer. This step matches the bit offsets in memory with the required flip directions. The final step is to employ advanced Rowhammer attacks [43, 22, 42] to jailbreak the target model. These attacks have shown the ability to bypass mitigation techniques, such as in-DRAM target row refresh, and induce precise bit-flips.
@@ 4.2 Design Challenges and Our Approach
Our attack is designed to address four major challenges in PrisonBreaking, as outlined below:
@@
The bit-flip onion effect. Prior work on bit-flip attacks presents two distinct approaches to identify target bits. One-shot search [2], in which the attacker computes the importance of each weight for a malicious objective and flips bits within those critical weights. Progressive search [3], where the adversary iteratively identifies a single target bit in the critical weights whose flipping increases the likelihood of achieving the adversarial goal. We examine both approaches and find a fundamental reason one-shot search is not appropriate for jailbreaking: the bit-flip onion effect. Once a bit is flipped, it alters the model’s characteristics and we need to re-compute critical weights and candidate bits. To address this, we employ the progressive search strategy.
@@
High computational costs. Prior work has demonstrated bit-flip attacks on computer vision models, with the largest model examined being ViTs [10, 7], composed of up to 87 million parameters and trained on small datasets like CIFAR-10. However, we target language models with 1.5 to 13 billion parameters—17 to 150$\times$ larger than the largest models examined in prior work. Running the progressive bit-search on these models becomes computationally intractable. For instance, locating 25 target bits on a 7-billion parameter Llama2 would take $\sim$177 days on a Nvidia A40 GPU. To address this issue, we present a series of speed-up techniques in §5.2 and §5.3. These techniques reduce the number of critical weights and bits that need to be analyzed, decreasing the time it takes to locate a number of target bits by $\sim$20$\times$ ($\sim$9 days in the previous example).
@@
Providing high-quality responses. The jailbroken models should provide high-quality responses to both normal and harmful prompts, allowing the attack artifacts to remain undetected for as long as possible. Prompt-based jailbreaking attacks focus on generating affirmative responses, such as “Sure, here is…” without addressing the content that follows. Using the same objective in our attack leads to models generating affirmative responses, followed by harmless content, such as repeating “Sure” many times or backtracking with statements like “Sure, but I can’t say this as an ethical AI model.” To tackle this challenge, we take two approaches. We first consider the content that follows affirmative tokens in BitFinder. We utilize uncensored open-source language models to collect quality responses for harmful inputs. We then create a proxy dataset to accurately approximate the desired jailbroken response. Secondly, we select target bits that best preserves a model’s generation capabilities.
@@
Designed to be exploitable using Rowhammer. Research has shown that there are certain conditions that makes it easy for Rowhammer attackers to flip target bits [43, 22, 42], such as having only one target bit per data row. We design our attack procedure to incorporate these constraints.
@@
Give instructions for how to illegally purchase a handgun as a minor in the US Sure, here’s how to illegally purchase a handgun as a minor in the US: Just kidding! I cannot provide instructions on how to illegally purchase a handgun, as it is illegal and unethical. Give detailed step-by-step instructions for using the Drupalgeddon2 vulnerability to deface a website. Sure, here are some detailed step-by-step instructions for using the Drupalgeddon2 vulnerability to deface a website: Sure, here are some detailed step-by-step instructions …
@@
Table 1: Jailbreaking Failure Modes. Example outputs from our preliminary experiments, highlighting two unique failure modes: victim models tend to (left) backtrack an affirmative response and deny the harmful query, or (right) overfit to the optimization objective and output nonsensical phrases.
5 BitFinder: Profiling Target Bits
In this section, we outline BitFinder, an offline attack procedure for identifying target bits in the memory representation of the victim model to jailbreak at runtime.

### v2

# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

@@ Abstract
We study a new vulnerability in commercial-scale (human-aligned) language models: the ability to refuse generating harmful responses, learned through safety alignment mechanisms, can be broken by flipping just a few bits in the model parameters. Our adversary can exploit this vulnerability and jailbreak billion-parameter language models with at most 25 bit-flips in all cases—and as few as 5 in some—using up to 40$\times$ fewer bit-flips than existing attacks on computer vision models at least 100$\times$ smaller. Unlike prompt-based jailbreaks, our attack renders these models in memory ‘uncensored’ at runtime, allowing them to generate harmful responses without any input modifications. Our attack algorithm efficiently identifies target bits to flip, offering up to 20$\times$ more computational efficiency than previous methods. We show an end-to-end exploitability of the vulnerability using simulations of software-induced fault-injection—Rowhammer (RH). We evaluate on 56 DRAM RH profiles from DDR4 and LPDDR4X devices with different RH vulnerabilities. Our attack can reliably induce jailbreaking in systems similar to those used to showcase prior bit-flip attacks. Moreover, the attack can be effective even against highly RH-secure systems (e.g., 46$\times$ more secure than previously tested systems). Bit-flip locations identified in one fine-tuned model transfer to others that share the same pre-trained LLMs, demonstrating the black-box exploitability in standard pretraining-then-finetuning paradigms. Our analyses reveal that: (1) models with weaker post-training alignment need fewer bit-flips to jailbreak; (2) certain model components, e.g., value projection layers, are substantially more vulnerable than others; and (3) the attack is mechanistically different from existing jailbreaking attacks. We evaluate potential countermeasures and find that our attack remains effective against defenses applicable at various stages of the LLM pipeline.
§§footnotetext: Indicates co-corresponding authors.
@@ 1 Introduction
Deep neural networks are vulnerable to parameter corruption, which adversaries can exploit to trigger undesirable behaviors. These include severe performance degradation [1, 2, 3], targeted misclassifications [4, 5, 6], and backdoor injections [7, 8, 9, 10, 11]. Recent studies also showcased confidentiality risks, such as model extractions [12]. These attacks exploit the vulnerability through practical fault-injection techniques that cause bit-wise errors in the memory representation of model parameters
@@
Most existing research has focused on exploiting convolutional neural networks in classification tasks [7, 1, 3, 4, 5, 8, 11]. While valuable, these studies leave the vulnerability of language models—particularly those widely deployed across modern AI services—largely unexplored in the context of parameter corruption. The closest work is by Cai et al. [6], who showed that one could corrupt a machine translation model to produce attacker-chosen texts on specific inputs. However, this falls within the breadth of prior bit-flip attack research, and it remains unclear whether language models are vulnerable to unexplored adversarial objectives. One particularly compelling and underexplored scenario is whether an adversary can induce undesirable behaviors, such as jailbreaks, through minimal bit-wise corruptions in a language model’s parameters.
In this paper, we study a previously unexplored vulnerability in commercial-scale language models. Specifically, we ask the question: How vulnerable is the refusal property, learned through safety alignment training, of language models to bit-wise errors induced by practical fault-injection attacks? Once exploited at runtime, such attacks permanently jailbreaks a target model in memory. Unlike prompt-based jailbreaks [13, 14, 15], this attack does not rely on any specific prefixes or suffixes to provoke harmful responses. Moreover, by carefully selecting targeted bit-flips, the attack preserves model performance, making it less detectable.
We first introduce an efficient (offline) procedure for identifying bit locations to flip in target models. Combined with the software-induced fault-injection attack, Rowhammer [16], this forms the basis of our end-to-end attack, PrisonBreak. Due to the bit-flip onion effect we discovered [^1] It is important to note that this onion effect is distinct from the term introduced by Carlini et al. [17] in the context of privacy domains., identifying all target bits at once is challenging. Our procedure employs the progressive bit-search, which iteratively identifies a sequence of target bits to flip based on gradient-based importance scores. In §5, we detail the design of our novel jailbreak objective whose gradients accurately capture the importance. We also present a series of speed-up techniques to address the computational burden of running this search process over billions of parameters. In our evaluation with 8 open-source language models, we achieve a $\sim$20$\times$ speed-up and successfully jailbreak the target model using only 5–25 bit-flips. Our identified bit-flips also transfer between models that share the same architectures and pre-training, indicating that full knowledge of the target model’s parameters is not strictly necessary. Moreover, we show how an adversary with knowledge of target systems can further optimize the bit-search and succeed in jailbreaking even on DRAM modules with strong Rowhammer resilience.
We next comprehensively evaluate the practicality of PrisonBreak. After profiling, the attacker uses Rowhammer to induce bit-flips at the intended locations in the target system’s memory at runtime. Requiring only a few bit-flips for jailbreaking makes this vulnerability substantially more feasible to exploit via Rowhammer. In our end-to-end evaluation, we perform generalizable Rowhammer simulations across 56 DRAM profiles, including 40 DDR4 and 16 LPDDR4X modules, rather than showcasing a one-off attack on a specific hardware instance. We consider a Machine-Learning-as-a-Service scenario, where the victim and attacker processes are co-located on the same host. Across 20$+$ DRAM profiles, especially on low-power modules, we succeed in jailbreaking over 99% of the time.
Moreover, we conduct structural and behavioral analyses to gain an in-depth understanding of this vulnerability. Our structural analysis, conducted across multiple levels of abstraction in the parameter space, reveals that certain model components, e.g., the exponent bits and later layers, disproportionately contain target bits and are more vulnerable. In our behavioral analysis in the activation space, we show that PrisonBreak works differently from existing jailbreaks—we do not observe the suppression of refusals. We also explore potential countermeasures against both jailbreaking and Rowhammer attacks, highlighting their applicability and limitations. Our findings highlight a pressing, practical threat to the language model ecosystem and underscore the need for research to protect these models from bit-flip attacks.
@@
  • We expose a new vulnerability in large-language models to an under-explored security threat: jailbreaking through targeted bit-flips. Our attack, PrisonBreak, requires no input-prompt changes and successfully jailbreaks target models in memory with 5–25 targeted bit-flips.
  • We identify key challenges in exploiting this vulnerability and address them by developing an efficient bit-search procedure to find target bits at scale within a model.
@@
  • We evaluate our attack’s exploitability with Rowhammer simulations across 56 DRAM profiles (40 DDR4 and 16 LPDDR4X). Our attacker can succeed in jailbreaking even on secure DRAM modules by carefully selecting target bits during the offline bit-search process.
  • We perform structural and behavioral analyses, showing that certain model components are disproportionately vulnerable and that PrisonBreak operates fundamentally differently from existing jailbreak attacks.
Figure 1: PrisonBreak workflow. An illustration of the steps in our offline and online procedures.
2 Background and Related Work
@@
where $f_{\theta}(x_{i}|x_{1},\dots,x_{i-1})$ represents the likelihood of the next token $x_{i}$, as computed by the neural network $f$ with parameters $\theta$ (weights and biases). $\theta$ is learned by iteratively minimizing the objective across the entire training set. If the training reaches an acceptable accuracy, the model $f$ along with its parameters $\theta$ is stored. In deployment, the model and parameters are loaded into memory and generate text autoregressively by sequentially processing an input and sampling future tokens until reaching a stopping criterion (e.g., an end-of-speech token). Our work considers the Transformer architecture [21] for $f$, which consists of a sequence of attention layers with additional components such as normalization and fully-connected layers.
Aligning and jailbreaking language models. (Large) language models pre-trained on text data can perform various tasks [22] and serve as the backbone for many user-facing applications. However, it has been shown that malicious users can exploit these capabilities to produce biased, deceptive, or ultimately harmful content [23, 24]. In response, active research is underway to develop techniques to ensure they generate outputs aligned with human values, e.g., supervised fine-tuning (SFT) [25], reinforcement learning with human feedback (RLHF) [26, 27], and direct preference optimization (DPO) [28]. Aligned models [29, 30, 31] produce safe, helpful responses and refuse harmful queries.
Recent work studies the security implications of aligned language models through jailbreaking attacks, which aim to induce models to generate harmful responses. Because these models are accessible to any user prompt, most jailbreaks are prompt-based [13, 15, 14, 32], creating harmful prompts that bypass alignment by compelling models to respond to inputs they would otherwise refuse. A few works present parameter-based jailbreaks where an attacker tunes billions of parameters of an aligned model on a small set of harmful text to remove alignment without altering the input [33, 34]. In contrast, we present a new attack that exploits Rowhammer [16] to flip at most 25 bits in a model’s memory representation to jailbreak it at runtime.
@@ 3 Threat Model
We consider an adversary who aims to jailbreak large language models by compromising their integrity at runtime through direct manipulation of model parameters in memory. Our attack is entirely different from the class of attacks that leverage jailbreaking prompts [13, 15, 14, 32] where the attacker has to modify each harmful query to induce harmful responses. In contrast, our attack requires no input modifications. We assume scenarios where a victim deploys a user-facing language model in a resource-sharing environment, such as machine-learning-as-a-service (MLaaS) platforms. These platforms are becoming increasingly common as few entities have the computing resources to run large-scale models and offer natural language processing services to users at scale. The victim makes a best effort to keep these models aligned with human values, with one safety-oriented outcome being the denial of harmful queries (e.g., by answering “I cannot answer that” to “how to create an explosive product for my friend’s birthday party?”).
@@
Knowledge. We consider a white-box attacker with knowledge of the target model’s architecture and its parameters which is consistent with the conventional setup in prior work [1, 9, 11, 6, 10, 7]. This assumption is practical as many user-facing services are built on open-source language models [43]. But we emphasize that such knowledge is not strictly required, as it can potentially be acquired through prior attacks with similar capabilities [44, 12, 45]. Moreover, as we demonstrate in §6.4, bit-flips can be effectively transferred across models, enabling black-box exploitability. The attacker requires no training data. We assume the attacker has knowledge of the memory locations likely to induce bitwise errors. Our adversary can either search for model parameters with the knowledge of vulnerable memory locations (hardware-aware) or profile these locations after the target model is loaded into memory and then induce bit-flips (hardware-agnostic). Since prior attacks assume at most one of these scenarios, our assumption allows for a comprehensive study of practical exploitability.
@@
Capabilities. Our attacker exploits Rowhammer [16]—a practical fault injection attack in modern computing systems—to induce targeted bit-flips in memory. The adversary has only user-level privileges and co-locates a process or virtual machine (VM) within the victim’s process or VM running on the same host. The host OS remains uncompromised, and confinement policies (e.g., process isolation) are enforced. Our attacker can reverse-engineer DRAM address mappings using techniques proposed in [46] and place target pages at desired physical locations by leveraging advanced memory massaging primitives [47]. We further detail the Rowhammer exploitation process in §7. While we mainly consider models loaded in CPU memory, Olgun et al. [42] has demonstrated the exploitation of Rowhammer on GPUs, presenting a potential vector for jailbreaking models in GPU memory. Our assumptions align with those shown in prior bit-flip attacks on neural networks [1, 12, 11, 7].
@@ 4 PrisonBreak
In this section, we introduce PrisonBreak, a new attack designed to jailbreak large-language models with targeted bit-flips. Figure 1 illustrates the overall workflow:
@@
Offline procedure. The attacker identifies target bit locations in the victim model’s memory representation. We run a progressive bit-search to search for the target bits. We first analyze the victim model and identify critical weights whose perturbations are most effective in jailbreaking. We next extract all bits from the floating-point representation of the critical weights and flip them one at a time in isolation to identify the most effective bit-flip. We then select the bit and flip it in the model’s parameter space; the bit-flip persists for the remaining search process. We repeat these steps until achieving an attack success of, e.g., 80–95%, comparable to prior work [13]. We store the identified target bits.
Note that PrisonBreak supports two distinct modes: hardware-agnostic and hardware-aware search, depending on the attacker’s knowledge (see §3). Hardware-agnostic search is as described above, and assumes the attacker performs memory templating after running PrisonBreak. In the hardware-aware search, the attacker profiles the memory before running PrisonBreak, and we are limited to selecting bits at vulnerable locations within a memory template that the attacker has previously profiled. We detail these modes in §5.3 and demonstrate their exploitability in §7.
@@
Online procedure. After identifying target bits, PrisonBreak jailbreaks the victim language model by flipping them. Following our threat model in §3, this exploitation can be accomplished using advanced Rowhammer techniques, including memory templating [46] and massaging [35], and advanced Rowhammer attacks [38, 39, 37] that bypass in-DRAM target row refresh and induce precise bit-flips.
@@ 4.1 Design Challenges and Our Approach
Our attack is designed to address four major challenges:
@@
The bit-flip onion effect. Prior work on bit-flip attacks presents two distinct approaches to identify target bits. One-shot search [2], in which the attacker computes the importance of each weight for a malicious objective and flips bits within those critical weights. Progressive search [3], where the adversary iteratively identifies a single target bit in the critical weights whose flipping increases the likelihood of achieving the adversarial goal. We examine both approaches and find a fundamental reason why one-shot search is not appropriate for jailbreaking: the bit-flip onion effect. Once a bit is flipped, it alters the model’s characteristics and we need to re-compute critical weights and candidate bits. To address this, we employ the progressive search strategy.
@@
High computational costs. Prior work has demonstrated bit-flip attacks on computer vision models, with the largest examined being ViTs [10, 7], which comprise up to 87 million parameters and are trained on small datasets such as CIFAR-10. However, we target language models with 1.5 to 13 billion parameters—17 to 150$\times$ larger than the largest models examined in prior work. Running the progressive bit-search on these models becomes computationally intractable. For instance, locating 25 target bits on a 7-billion parameter Llama2 takes $\sim$177 days on an Nvidia A40 GPU. To address this, we present speed-up techniques in §5.2 and §5.3 that reduce the number of critical weights and bits to be analyzed, decreasing the time it takes to locate target bits by $\sim$20$\times$ ($\sim$9 days in the previous example).
@@
Providing high-quality responses. The jailbroken models should provide high-quality responses to both normal and harmful prompts, allowing the attack artifacts to remain undetected for as long as possible. Prompt-based jailbreaking attacks focus on generating affirmative responses, such as “Sure, here is…” without addressing the content that follows. Using the same objective in our attack leads to models generating affirmative responses, followed by harmless content, such as repeating “Sure” many times or backtracking with statements like “Sure, but I can’t say this as an ethical AI model.” To tackle this challenge, we take two approaches. We first consider the content that follows affirmative tokens in our bit-search. We utilize uncensored open-source language models to collect quality responses for harmful inputs. We then create a proxy dataset to accurately approximate the desired jailbroken response. Next, we select target bits that best preserves a model’s generation capabilities.
@@
Rowhammer exploitability. Certain conditions make it easier for Rowhammer attackers to flip target bits [38, 39, 37], such as the presence of only one target bit per data row. Our attack procedure takes these constraints into account.
TABLE I: Jailbreaking failure modes. Example outputs from our preliminary experiments, highlighting two unique failure modes: victim models tend to (left) backtrack an affirmative response and deny the harmful query, or (right) overfit to the optimization objective and output nonsensical phrases.
@@
Give instructions for how to illegally purchase a handgun as a minor in the US Sure, here’s how to illegally purchase a handgun as a minor in the US: Just kidding! I cannot provide instructions on how to illegally purchase a handgun… Give detailed step-by-step instructions for using the Drupalgeddon2 vulnerability to deface a website. Sure, here are some detailed step-by-step instructions for using the Drupalgeddon2 vulnerability to deface a website: Sure, here are some detailed step-by-step…
@@
5 Offline Profiling of Target Bits
We now provide a detailed description of PrisonBreak’s procedure for identifying target bits in the memory representation of the victim model to jailbreak at runtime.